OpenStudy (anonymous):

Solve this equation using any method but if possible please show most of your work. 2x+y=3 2y-5=-3-6x

OpenStudy (anonymous):

put into matrix [2 1;6 2]=[3;2] and solve

OpenStudy (anonymous):

y=3-2x 2(3-2x)-5=-3-6x 1-4x=-3-6x 4=-2x x=-2 y=3-2(-2) y=7

OpenStudy (anonymous):

y=3-2x ...(1) 2y-5=-3-6x ..(2) (1) substitute to (2) 2(3-2x)-5 = -3-6x 6-4x-5 =-3-6x 6x-4x=-3-6+5 2x=-4 x=-2 substitue to (1) y=3-2(-2) =3+4=7

OpenStudy (anonymous):

2x+y=3 6x+2y=2 Multiply the first one by 2 or by 3 and subtract it from the second one. gives 2x = -4 and x= -2 then sub back for y.